OpenStudy (anonymous):

x^2-44+k=0, solve for k

OpenStudy (cwrw238):

just subtract (x^2 - 44) from both sides

OpenStudy (anonymous):

so will it be 0+44-x^-k

OpenStudy (anonymous):

-x^2-k+44

OpenStudy (anonymous):

confused

OpenStudy (cwrw238):

x^2-44+k=0 k = -(x^2 - 44) k = 44 - x^2

OpenStudy (anonymous):

my choices are 16, 28, -28, 60, -60

OpenStudy (anonymous):

hmmmm...how do i solve this, 44 cant be squared

OpenStudy (tester97):

what you can do is plug in all the answers then if it equals 0 then that is the answer i no it takes time but thats what i would recommend if your are confused

OpenStudy (tester97):

*plug in the answers for k

OpenStudy (anonymous):

thanks, that helps alot :) cool

OpenStudy (tester97):

no prob i used this method a lot in algebra

OpenStudy (anonymous):

im so use to solving for everything and not having the option of multiple choice, i didnt think of that

OpenStudy (tester97):

but this only works for multiple choice when it is short answer you would have to get some help there

OpenStudy (cwrw238):

assuming x is a whole number then k = 28 28 = 44 - 4^2

OpenStudy (anonymous):

thanks! that was right

OpenStudy (cwrw238):

yw